1. Define economics, distiguish between microeconomics and macroeconomics and explain the basic economics problems.
  2. Ditinguish between changes in quantity demanded and changes in demand.
  3. How is a shortage different from a surplus?
  4. People read that drinking orange juice helps prevent heart disease. What is the effect on the equilibrium price and quantity of orange juice?
  5. The cost of memory chips used in computers falls. What is the effect on the equilibrium price and quantity of computers?
  6. You are told that a 10% increase in the price of a good has led to a 1% increase in the quantity supplied of the good after one month. Use the information to answer the following questions?

a. How you decribe the supply of this good?

b. Calculate the price elasticity of supply.

c. If after one year, the quantity supplied has inceased by 25%, calculate the new elasticity of supply.



1
Expert's answer
2022-01-26T10:00:30-0500

Solution:

1.). Economics is the study of how society uses its limited resources. Economics is a social science that studies the production, distribution, and consumption of goods and services.

Macroeconomics is the branch of economics that studies the overall operation of a country's economy. It is more concerned with the big picture and the analysis of factors such as growth, inflation, interest rates, unemployment, and taxes.

Microeconomics is the branch of economics that studies how households and businesses make purchasing, saving, price setting, and business competition decisions.

Microeconomics is concerned with individual decisions, whereas macroeconomics is concerned with decisions that affect entire countries and society as a whole.

The basic economic problems are what to produce, how to produce, for whom to produce, and what provisions are to be made for economic growth.

 

2.). When demand changes, the entire demand curve shifts to the left or right. A change in quantity demanded is a movement along the demand curve caused solely by a price change. The demand curve does not move in this case; rather, we move along the existing demand curve.

 

3.). Surplus refers to the amount of a resource that exceeds the amount actively used, which occurs when prices or quantities are excessively high. A shortage, on the other hand, refers to a condition in which there is an excess demand for products in comparison to the quantity supplied in the market, i.e. when there is insufficient supply of a product.

 

4.). The equilibrium price and quantity of orange juice will increase due to the increased demand for orange juice. The increased demand will push the orange juice price higher, and suppliers will supply more orange juice due to the increased margins.

 

5.). The equilibrium price of computers will decrease while their quantity will increase. Computers will be cheaper, and they will be able to be supplied at lower prices in the market.

 

6.). a.). The supply of this good is price inelastic. It has low responsiveness to price changes.

 

b.). Price elasticity of supply = "\frac{\%\;change\; in\; quantity\; supplied}{\%\; change\; in\; price}"

PES = "\frac{1\%}{10\%} = 0.1"

 

c.). Price elasticity of supply = "\frac{\%\;change\; in\; quantity\; supplied}{\%\; change\; in\; price}"

PES = "\frac{25\%}{10\%} = 2.5"